OpenStudy (anonymous):

A.) A woman spent $250 for jeans and a ski jacket. If she spent y dollars for the ski jacket, represent the amount she spent for the jeans. B.) The width of a rectangle is x centimeters. Represent the lenghth of the rectangle if it exceeds twice the width by 3 centimeters

OpenStudy (anonymous):

Since the sum of the two values must equal 250, then we have: \[ j+y=250 \]Solving for \(j\) we get: \[ j=250-y \]Since: \[ (250-y)+y=250 \]Then, the amount spent on jeans is: \[ 250-y \]And so, we're done.

OpenStudy (anonymous):

I dont really get how you did that

OpenStudy (anonymous):

i feel so stupid

OpenStudy (anonymous):

i get it up to Since: (250−y)+y=250 Then, the amount spent on jeans is: 250−y And so, we're done.

OpenStudy (anonymous):

It's all right, here: So we know we have the two values, which must add up to 250. Using that, we get: \[ j+y=250 \]Subtract y from both sides: \[ j+y-y=250-y\implies\\ j+0=250-y\implies\\ j=250-y \]So, the price for the jeans is \(250-y\), where \(y\) is the price for the jacket.

OpenStudy (anonymous):

Oh, that's just to show that the expressions are equal, it's an added detail, not really all that important.

OpenStudy (anonymous):

thanks so muchh!!!!!!!!! can you also help me with b? its fine if you cant but thanks again

OpenStudy (anonymous):

Sure, sorry I was out eating. If the length exceeds twice the width (which we call \(x\)) by three, then: \[ 2x+3=l \]

OpenStudy (anonymous):

thanks thats what i got

OpenStudy (anonymous):

Nice, and sure thing.